OpenStudy (anonymous):

4x-1=7 solve equation,and please explain step by step?

OpenStudy (anonymous):

4x-1=7 4x=7+1 4x=8 x=8/4 x=2

OpenStudy (anonymous):

ah,so i was right..stupid book.lol thanks =)

OpenStudy (anonymous):

anytym ! :-)